# Generative AI for Cel-Animation: A Survey

### v4

@@ 3.3 GenAI for Post-production
Compositing requires more than simply overlaying foregrounds on backgrounds; it must ensure consistency in lighting, perspective, and spatial coherence. While layout unifies general geometry, the photography phase handles nuanced refinements, such as matching foreground lighting to scene ambience, adding shadows, and integrating particle or texture effects. Studies [24, 25, 35, 48, 113, 75] explore background-foreground harmonization. For instance, ObjectStitch [103] refines foreground placements for seamless integration, while ICLight [3] enables editable, consistent lighting. These techniques collectively advance realism and stylistic coherence in final renders.
@@
  • Music Generation: Music in animation often serves to amplify character emotions and enhance the atmosphere. For instance, an energetic song can heighten the intensity of a battlefield scene, while a melancholic tune can evoke empathy and draw the audience into a poignant narrative. AI-driven music generation primarily leverages text-guided methods, including diffusion models [67, 69, 43] and sequence-to-sequence modeling [2, 26, 44]. However, precise synchronization with visual content requires specialized video-guided models such as FoleyMusic [31], Video2Music [51], V2Meow [106], MelFusion [20], and VidMuse [114]. Although not specifically tailored for anime, these methods offer critical foundations for synchronization in AI-driven animation.
  • Sound Effects Generation: Effective sound effects require precise alignment with visual cues, encompassing various scene-specific noises (e.g., animal calls, mechanical sounds). Recent transformer-based models (SpecVQGAN [46], Im2Wav [89], FoleyGen [76]) and diffusion-based approaches [74, 12, 63] address semantic and temporal alignment, achieving more realistic and synchronized sound effects.
@@
DB synchronizes dialogue audio with character lip movements and expressions, traditionally performed through either pre-animated adjustments or post-animation recording. GenAI approaches, notably EmotiVoice [23], automate this by detecting facial regions and adjusting or generating facial movements guided by reference audio, subtitles, and visual context. Such models fuse acoustic features, phoneme-level precision, and emotional cues to ensure natural and expressive audio-visual synchronization, significantly streamlining the DB process.
@@
Currently, no dedicated GenAI methods specifically address animation cutting. However, existing general-purpose approaches like Reframe Anything (RAVA) [8], which uses LLM-generated FFmpeg instructions for reframing, and M-SAN [109], which extracts coherent segments from longer videos, show potential for adaptation to animation cutting. These methods can aid in efficiently managing duration and preserving content saliency.
@@ 3.4 Others
Context-aware editing methods enhance precision and controllability in cel-animation modifications. Re:Draw [9] blends inpainting and translation methods to ensure local edits remain contextually coherent and visually consistent, utilizing dual discriminators and specialized adversarial losses. Similarly, ScalingConcept [42], a training-free editing technique based on pre-trained diffusion models, addresses common post-production issues such as blurred lines and degraded clarity, significantly improving animation visual quality without compromising artistic style.
@@
Cel decomposition into reusable sprite layers facilitates effective editing and analysis. Sprite-from-Sprite [145] introduces a self-supervised framework handling varying complexity of animated elements. Additionally, Generative Omnimatte [57] employs generative video priors to robustly handle occlusions, dynamic backgrounds, and environmental effects, thus preserving details like shadows and reflections, further enriching decomposition capabilities.
@@
AI-assisted 3D methods provide robust tools for enhancing cel-animation dynamics. DrawingSpinUp [153] enables automatic 3D animation from character sketches, maintaining artistic integrity while supporting complex movements. Toon3D [124] reconstructs cartoon drawings into navigable 3D scenes, allowing novel viewpoints and sophisticated sequence planning. These tools bridge traditional 2D animation with 3D capabilities, offering greater flexibility and richer visual storytelling.
@@ 4 Case Study
Rock, Paper, Scissors [27] (2023).

### v5

@@ 3.3 GenAI for Post-production
Compositing requires more than simply overlaying foregrounds on backgrounds; it must ensure consistency in lighting, perspective, and spatial coherence. While layout unifies general geometry, the photography phase handles nuanced refinements, such as matching foreground lighting to scene ambience, adding shadows, and integrating particle or texture effects. Studies [cong2020dovenet, cong2022high, guerreiro2023pct, jiang2021ssh, tarres2024thinking, man2024floating] explore background-foreground harmonization. For instance, ObjectStitch [Song2023objectstitch] refines foreground placements for seamless integration, while ICLight [zhang2024iclight] enables editable, consistent lighting. These techniques collectively advance realism and stylistic coherence in final renders.
@@
  • Music Generation: Music in animation often serves to amplify character emotions and enhance the atmosphere. For instance, an energetic song can heighten the intensity of a battlefield scene, while a melancholic tune can evoke empathy and draw the audience into a poignant narrative. AI-driven music generation primarily leverages text-guided methods, including diffusion models [liu2023audioldm, audioldm2-2024taslp, huang2023noise2music] and sequence-to-sequence modeling [agostinelli2023musiclm, copet2024simple, huang2024audiogpt]. However, precise synchronization with visual content requires specialized video-guided models such as FoleyMusic [FoleyMusic2020], Video2Music [KANG2024123640], V2Meow [su2024v2meow], MelFusion [chowdhury2023melfusion], and VidMuse [tian2024vidmuse]. Although not specifically tailored for anime, these methods offer critical foundations for synchronization in AI-driven animation.
  • Sound Effects Generation: Effective sound effects require precise alignment with visual cues, encompassing various scene-specific noises (e.g., animal calls, mechanical sounds). Recent transformer-based models (SpecVQGAN [SpecVQGAN_Iashin_2021], Im2Wav [sheffer2022i], FoleyGen [mei2024foleygen]) and diffusion-based approaches [luo2024diff, chen2024action2sound, liang2024language] address semantic and temporal alignment, achieving more realistic and synchronized sound effects.
@@
DB synchronizes dialogue audio with character lip movements and expressions, traditionally performed through either pre-animated adjustments or post-animation recording. GenAI approaches, notably EmotiVoice [cong2024styledubber], automate this by detecting facial regions and adjusting or generating facial movements guided by reference audio, subtitles, and visual context. Such models fuse acoustic features, phoneme-level precision, and emotional cues to ensure natural and expressive audio-visual synchronization, significantly streamlining the DB process.
@@
Currently, no dedicated GenAI methods specifically address animation cutting. However, existing general-purpose approaches like Reframe Anything (RAVA) [cao2024reframe], which uses LLM-generated FFmpeg instructions for reframing, and M-SAN [Tang_2022_ACCV], which extracts coherent segments from longer videos, show potential for adaptation to animation cutting. These methods can aid in efficiently managing duration and preserving content saliency.
@@ 3.4 Others
Context-aware editing methods enhance precision and controllability in cel-animation modifications. Re:Draw [cardoso2024re] blends inpainting and translation methods to ensure local edits remain contextually coherent and visually consistent, utilizing dual discriminators and specialized adversarial losses. Similarly, ScalingConcept [huang2024scaling], a training-free editing technique based on pre-trained diffusion models, addresses common post-production issues such as blurred lines and degraded clarity, significantly improving animation visual quality without compromising artistic style.
@@
Cel decomposition into reusable sprite layers facilitates effective editing and analysis. Sprite-from-Sprite [zhang2022sprite] introduces a self-supervised framework handling varying complexity of animated elements. Additionally, Generative Omnimatte [lee2024generativeOmnimatte] employs generative video priors to robustly handle occlusions, dynamic backgrounds, and environmental effects, thus preserving details like shadows and reflections, further enriching decomposition capabilities.
@@
AI-assisted 3D methods provide robust tools for enhancing cel-animation dynamics. DrawingSpinUp [zhou2024drawingspinup] enables automatic 3D animation from character sketches, maintaining artistic integrity while supporting complex movements. Toon3D [weber2024toon3d] reconstructs cartoon drawings into navigable 3D scenes, allowing novel viewpoints and sophisticated sequence planning. These tools bridge traditional 2D animation with 3D capabilities, offering greater flexibility and richer visual storytelling.
@@ 4 Case Study
Rock, Paper, Scissors [corridorRPS2023] (2023).